Make a table for interior angles of regular polygan having 5 sides to 10 sides.
Derive a general formula to find the interior angle of a regular n-sided polygon.

Answer by Alan3354(69443)

Make a table for interior angles of regular polygon having 5 sides to 10 sides.
Derive a general formula to find the interior angle of a regular n-sided polygon.
---------------
Make the formula first:
Int angles = 180*(n-2)/n where n = # of sides
------
Then make a table.
=========================
Another method:
Ext angles = 360/n
Int angle = 180 - Ext angle